Rewrite the equation 4x-6y=-30 as a function of x.
Bob invested $20,000, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%?
One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
The lenght of a rectangle is 8 in., more than twice its width. If the perimeter of the rectangle is 64., find the width of the rectangle.
Thank you for helping me out. I really appreciate it.



Answer by stanbon(75887)

Rewrite the equation 4x-6y=-30 as a function of x.
6y=4x+30
3y=2x+15
y=(2/3)x+5
----------------------
Bob invested $20,000, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%?
Let amt. invested at 14% be "x"; interest on this is 0.14x dollars
Amt. invested at 13% will be "20,000-x"; interest on this is 0.13(20000-x)
= 2600-0.13x dollars
EQUATION:
interest + interest = 2720
0.14x + 2600-0.13x = 2720
0.01x = 120
x = $12,000 (invested at 14%)
20000-x = $8,000 (invested at 13%)
-------------------
One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Let one number be "x"; The other number is "x+12"
EQUATION:
x+2(x+12)=39
3x+24=39
3x = 15
x=5
------------------
The length of a rectangle is 8 in., more than twice its width. If the perimeter of the rectangle is 64., find the width of the rectangle.
Let width = x inches
Length = 2x+8 inches
Perimeter = 2(length+width)
64 = 2(2x+8+x)
32 = 3x+8
3x=24
x=8 inches
width = 8 inches
==============
Cheers,
Stan H.